A meter stick is held perpendicular to the ground. It casts a shadow 1.5 meter long. At the same time, a telephone pole casts a shadow that is 9 meters long. How tall is the telephone pole?
Found 2 solutions by scott8148, richwmiller:
Answer by scott8148(6628)

the shadow of the stick is one and a half times as long as the stick

so the shadow of the pole is also one and a half times as long as the pole

9 / (3/2) = 6

Answer by richwmiller(17219)

stick/shadow and pole/shadow
1/1.5=p/9
p=6
or
1.5/9=1/x
9=1.5x
6=x